OpenStudy (anonymous):

Solve on the Interval [0, 2pi); (sin x-1)(2 sin^2 x-5 sin x + 2)=0

OpenStudy (anonymous):

A. x= pi/2, x= pi/6, x=5pi/6 B x= pi, x= 2pi/3, x= 5pi/3 C. x=2pi, x= pi/2, x=5pi/4 D.x= 2pi, x= pi/2, x=pi/3

OpenStudy (anonymous):

A